### v3

# Puzzle Solving using Reasoning of Large Language Models: A Survey

Abstract
Exploring the capabilities of Large Language Models (LLMs) in puzzle solving unveils critical insights into their potential and challenges in AI, marking a significant step towards understanding their applicability in complex reasoning tasks. This survey leverages a unique taxonomy—dividing puzzles into rule-based and rule-less categories—to critically assess LLMs through various methodologies, including prompting techniques, neuro-symbolic approaches, and fine-tuning. Through a critical review of relevant datasets and benchmarks, we assess LLMs’ performance, identifying significant challenges in complex puzzle scenarios. Our findings highlight the disparity between LLM capabilities and human-like reasoning, particularly in those requiring advanced logical inference. The survey underscores the necessity for novel strategies and richer datasets to advance LLMs’ puzzle-solving proficiency and contribute to AI’s logical reasoning and creative problem-solving advancements.
Puzzle Solving using Reasoning of Large Language Models: A Survey
Panagiotis Giadikiaroglou, Maria Lymperaiou, Giorgos Filandrianos, Giorgos Stamou Artificial Intelligence and Learning Systems Laboratory School of Electrical and Computer Engineering National Technical University of Athens panosgiadi@gmail.com, {marialymp, geofila}@islab.ntua.gr, gstam@cs.ntua.gr
1 Introduction
Recent developments in LLMs such as GPT-3 Brown et al. (2020) and GPT-4 OpenAI et al. (2023) have showcased their logical reasoning abilities across various domains Liu et al. (2023a, b); Bao et al. (2023); Creswell et al. (2022). Despite these advances and their demonstrated capabilities in deductive reasoning Saparov et al. (2023), LLMs face limitations in inductive reasoning settings, as analyzed by Xu et al. (2023a); Bang et al. (2023). The specific application of LLMs to puzzle solving, has not been thoroughly summarized.
Figure 1: Riddle from RiddleSense Lin et al. (2021). GPT-4, LLaMA2-70B and Bard chose the right answer.
Our main contributions are as follows: 1 We introduce a distinction between rule-based and rule-less puzzles (§2), highlighting the varied knowledge demands necessary to tackle them. 2 We analyze the methodologies LLMs use to solve puzzles (§3), assessing their impact on each category and comparing them with conventional problem-solving methods. 3 A detailed exploration of existing benchmarks that gauge models’ reasoning abilities is conducted (§4). 4 Finally, this paper offers a detailed view of the present obstacles faced in puzzle-solving with LLMs and highlights a wide array of prospects for future research (§5).
Our categorization diverges from existing logical reasoning taxonomies by emphasizing on the underlying cognitive processes and the skills required for puzzle solving, rather than the question format Luo et al. (2023) or the nature of reasoning (deductive, inductive, abductive) Luo et al. (2023); Yu et al. (2023a); Yang et al. (2023b); Qiao et al. (2022); Huang and Chang (2022); Flach and Kakas (2000). For instance, the existence of rules in puzzles such as Sudoku, Crosswords, or Minesweeper necessitates additional skills (e.g. strategy development) to correctly understand the game’s rules or the ability to correctly format the output. In contrast, rule-less puzzles, such as riddles (Figure 1), programming challenges, and commonsense reasoning problems, leverage the model’s inherent knowledge for solution derivation.
In our work, we define puzzles as problems that test cognitive abilities including logical reasoning, spatial cognition, and creative thinking by requiring the solver to discern patterns, apply deduction, and combine insights from available information in order to arrive at the correct solution. Notably, we exclude puzzles that cannot be expressed in text in any way, such as jigsaw puzzles Markaki and Panagiotakis (2022), or problems that require multimodal understanding abilities of LLMs Chia et al. (2024); Ghosal et al. (2024). Mathematical puzzles are also excluded, as this area diligently covered by the recent work of Liu et al. (2023c).
We keep track of the latest progress in the field of puzzle solving using LLM reasoning in our GitHub https://puzzlellms.github.io/.
{forest}
[…]
Figure 2: A taxonomy of Puzzle Categories with the corresponding Datasets.
2 Categorization of Puzzle Problems
In assessing LLMs’ reasoning capabilities, it is essential to categorize puzzles into coherent groups. We distinguish puzzles by their reliance on formal rules or broader world knowledge accompanied by general inferential skills, as illustrated in Figure 2. This categorization not only highlights the cognitive diversity puzzles present, but also aligns with distinct reasoning challenges: rule-based puzzles demand logical deduction and strategic foresight within closed environments with defined parameters, whereas rule-less puzzles require general reasoning abilities, interpreting situations and explaining events by drawing inferences based on practical knowledge about the everyday world.
By separating puzzles into these categories, we aim to provide a nuanced analysis of LLMs’ problem-solving abilities, reflecting on both structured challenges and those necessitating broader inferential reasoning.
2.1 Rule-based Puzzles
Rule-based Puzzles provide the model with explicit victory conditions, legal move sets or state transition rules. We further subdivide this category based on whether the state transitions are deterministic or incorporate randomness.
Deterministic games always produce the same successor state given a current game state and action taken according to the rules. For example, in Chess, making a move always yields one unambiguous new board layout. Other examples include Sudoku, maze navigation, or solving a Rubik’s cube. The model should learn strategies that operate within the possibility space defined by legal game mechanics.
Stochastic games incorporate randomness or hidden information, i.e. the same player action can lead to different probability distributions over next states. Examples include Minesweeper (hidden bomb locations) or card games e.g. Poker where opponents hold private hands. Mastering these games requires reasoning over uncertain states, planning multiple moves in advance and managing risk.
Thus, while both subgroups require logical reasoning bounded by formal rules, stochastic games pose the additional challenge of decision-making under uncertainty. Excelling in deterministic games enables pure reliance on deduction and forward search, while stochastic environments also require abilities for probabilistic inference, risk analysis, and reasoning with incomplete information.
2.2 Rule-less Puzzles
Unlike rule-bounded puzzles, rule-less problems rely more on flexible thinking and real-world knowledge to interpret vague situations and infer unobserved details. Rather than testing systematic search or strategic planning, these puzzles measure cognitive skills for contextual interpretation, conceptual combination, and reasoning from common experiences. The following fall under this category.
Riddles utilize clever wordplay and literary devices to conceal answers. For example, "What gets wetter the more it dries?" obscures the solution of "a towel" through metaphor. Solving riddles requires making abstract connections between concepts hidden in lyrical language. This assesses skills for fluid reasoning, conceptual blending, and lateral thinking to decode linguistic relationships.
Programming Puzzles provide code snippets and require analyzing or modifying the underlying program logic. Schuster et al. (2021) define a programming puzzle as a short Python program $f$, and the goal is to find an input which makes $f$ return True. Such puzzles assess skills like tracing execution, fixing errors, or anticipating outputs based on coding semantics. For example, the following puzzle tests understanding programming semantics to predict a system’s behaviour:
def mystery(x):
return x // 2
print(mystery(10))
Commonsense Reasoning Puzzles depict typical situations omitting key details. Solvers must explain events by inferring plausible implicit assumptions about motivations, causes and effects. For instance, the question "A man who was outside in the rain without an umbrella or hat didn’t get a single hair on his head wet. Why?" requires pragmatic analysis of unstated contextual factors.
3 Methods and Strategies
In applying LLMs to puzzle solving, a wide array of methods and strategies enhances complex reasoning and performance. This section outlines the approaches used to address puzzles, aiming to highlight their application within this unique context. Given the extensive literature on prompt engineering and related methods Besta et al. (2024); Chen et al. (2023); Yu et al. (2023b); Chu et al. (2023); Qiao et al. (2022); Liu et al. (2021), we concentrate on the techniques most prevalent for puzzle solving, instead of describing each method separately. We divide existing methods into prompting techniques, neuro-symbolic approaches for puzzle translation and fine-tuning for specific domains. A detailed overview of the methods utilized across different puzzle categories is presented in Table 1. We also discuss how conventional methods have faced these problems before the LLM era (App. A.2).
3.1 Prompting Methods
Prompting strategies that provide intermediate reasoning steps are pivotal in enhancing the puzzle-solving capabilities of language models. The few-shot in-context learning paradigm offers one or more demonstrations within prompts, significantly improving performance for both rule-based and rule-less puzzles by showcasing the reasoning process without additional training Brown et al. (2020); Dong et al. (2023); Zhou et al. (2022).
Recent works focus on how different ‘thought structures’ can guide LLMs to the final solution Besta et al. (2024).
Chain topologies, which include Chain-of-Thought (CoT) Wei et al. (2022); Kojima et al. (2022) have been applied to all kinds of puzzles, demonstrating their superiority over simple IO prompts. Self-Refine Madaan et al. (2023) is used for the Game of 24 (rule-based/deterministic), outperforming CoT with a 13% higher success rate Yao et al. (2023). Gu et al. (2023) use several methods in a rule-less detective-style benchmark, including Automatic CoT, which autonomously generates diverse reasoning chains for various questions Zhang et al. (2022); Complexity CoT leverages the complexity of prompted chains, where more intricate reasoning steps often lead to improved performance in complex inference tasks by selecting outcomes that demonstrate deeper reasoning capabilities Fu et al. (2022); and the Plan-and-Solve (PS) method, which uses two prompts for each problem—one to generate the reasoning process and the corresponding answer, and another to extract the final answer from the initial generation Wang et al. (2023a). Despite the varied approaches, none of these methods clearly outperformed CoT across all tested LLMs. The best results are achieved by Detective Thinking Prompt, a CoT-like method introduced in the same study, which does not exceed the 61.6% accuracy score of the best model, GPT-4. The method encourages the model to consider and analyze multiple clues within a given scenario, sequentially building towards a conclusion. This type of prompting can help the model handle complex scenarios where synthesizing disparate information correctly is crucial to generating accurate and logical outcomes. Schuster et al. (2021) exclusively utilized the solutions to programming puzzles that the model had already solved as examples, surpassing alternative approaches.
Tree topologies cover a variety of methods. Self-Consistency (SC) Wang et al. (2022) has been tested on rule-based/deterministic puzzles, such as the 8-puzzle, Game of 24 and Pocket Cube, as well as on rule-less commonsense reasoning puzzles, showcasing a small gain in the first category over CoT Ding et al. (2023); Yao et al. (2023); Mo and Xin (2023) and no clear benefit in the second one Gu et al. (2023). Tree-of-Thought(s) (ToT) Yao et al. (2023); Long (2023) has been exclusively applied to rule-based/deterministic puzzles so far, achieving significantly improved success rates over CoT, with increases ranging from 26% Mo and Xin (2023) to 70% Yao et al. (2023) depending on the puzzle and the depth of the tree, despite the increased LLM invocations Ding et al. (2023). Tree-of-Uncertain-Thought (TouT) Mo and Xin (2023) achieved even better results than ToT on the same challenges, with a 9% higher success rate on the Game of 24 and 3% on mini-crosswords. Finally, Inference-Exclusion-Prompting (IEP) Tong et al. (2023) employs a combination of forward and backward reasoning to approximate human logic and delivered some of the best results on riddles and commonsense puzzles when combined with CoT, scoring 82% on puzzles–up from 81% with zero-shot CoT–and 79% on riddles, compared to 82% with zero-shot CoT.
Graph topologies entail the following: Graph-of-Thought(s) (GoT) Besta et al. (2023); Lei et al. (2023) and Everything-of-Thought (XoT) Ding et al. (2023) have been used to solve rule-based/deterministic puzzles. While GoT has shown poorer results compared to ToT, with a decrease ranging from 2% to 6% Ding et al. (2023), XoT has been recognized as the most effective method for these puzzles. XoT integrates Monte Carlo Tree Search (MCTS) with LLMs for enhanced thought generation, achieving improvements in results from 53% to 69% compared to ToT. Additionally, XoT presents the fewest LLM invocations among the methods tested, including CoT, SC, ToT, and GoT.
A brief analysis of some basic methods not described here is presented in Appendix A.1, while a more detailed analysis of all the methods discussed can be found in the extensive work of Besta et al. (2024). Beyond the aforementioned methods, the use of extra information such as hints for riddles and commonsense puzzles, or introductions and summarizations of the puzzles, has also been employed. The inclusion of supplementary details appears to yield positive results, although this is not always the case; for instance, Chinese riddles typically show worse results when hints are used Zhang and Wan (2021).
3.2 Puzzle Translation
In this subsection, we summarize the neuro-symbolic techniques used by LLMs to translate text puzzles from natural language into forms more amenable to solutions by external tools. Notably, these methods do not test the LLMs’ puzzle solving capacity but rather assess their ability to encode puzzles into appropriate representations.
The primary approach involves using LLMs to generate logic rules from the puzzle’s natural language and subsequently solve it using a symbolic solver. Ishay et al. (2023) employ GPT-3 and GPT-4 to transform logic puzzles, such as chess puzzles, Jobs puzzle and Sudoku (rule-based/deterministic) into Answer Set Programming (ASP) formats by generating predicates and rules. They demonstrate that this method achieved significant results, with GPT-4 scoring 92% accuracy in a logic puzzles dataset Mitra and Baral (2015), compared to 7% in few-shot and 21% in zero-shot settings with the same model. They note that in few-shot settings, LLMs can generate complex programs that humans can easily refine and correct in case of code errors. Additionally, similar frameworks such as Logic-LM Pan et al. (2023a), LINC Olausson et al. (2023) and Yang et al. (2023a)’s method show promising results in logical reasoning tasks, although not specifically in puzzle settings.
While neuro-symbolic approaches have been applied to puzzle translation into logic rules, we have found no studies on transforming puzzles from natural language into code. However, techniques such as Program of Thoughts (PoT) prompting Chen et al. (2022) and Program-Aided Language (PAL) Gao et al. (2022) employ models to convert reasoning into Python programs for logical and mathematical reasoning datasets. Therefore, we encourage the research community to explore these methods for puzzle-solving tasks as well.
Given the structured nature of rule-based puzzles, this approach is inherently suitable for them. Consequently, it is logical that no studies have yet been conducted on rule-less puzzles in this context.
3.3 Fine-Tuning
Fine-tuning LLMs emerges as a potent strategy for enhancing their reasoning capabilities, ranging from general logical reasoning to specific puzzle-solving skills.
Logical Reasoning
LoGiPT Feng et al. (2023a) is a language model fine-tuned to excel in logical reasoning tasks by mimicking the symbolic reasoning of logical solvers. The fine-tuning process involves constructing an instruction-tuning dataset comprising natural language (NL) logical questions paired with symbolic reasoning steps. It is fine-tuned to bypass syntax errors typically encountered in NL to symbolic language parsing, enabling it to directly produce answers. LogiT5 Luo et al. (2023) leverages a multi-task learning approach, incorporating diverse datasets to enhance its reasoning capabilities across different logical domains. The model is fine-tuned on the LOGIGLUE benchmark, which includes various logical reasoning datasets, enabling it to perform better on tasks with limited data by transferring knowledge across tasks.
Rule-based Puzzles
In the domain of rule-based deterministic puzzles, Noever and Burdick (2021) observe suboptimal results when fine-tuning GPT-2 on Sudoku, Rubik’s Cube and Mazes, potentially due to a brief fine-tuning period and limited training examples. Regarding crosswords, various studies Rozner et al. (2021); Efrat et al. (2021) show mixed results, with some fine-tuned LLMs outperforming non-neural baselines and others not, highlighting the inherent challenge of cryptic crosswords for LLMs. Kazemi et al. (2023) demonstrate that fine-tuning LLMs with proofs and CoT under rule-based contexts yields some of the best results.
Rule-less Puzzles
In the realm of riddles, the study of Lin et al. (2021) illustrates that models like BERT Devlin et al. (2019), RoBERTa Liu et al. (2019) and ALBERT Lan et al. (2019) perform better when trained on both RiddleSense Lin et al. (2021) and CommonsenseQA Talmor et al. (2019) datasets, leveraging commonsense knowledge effectively. Moreover, Zhang and Wan (2021) report that combining fine-tuning on ALBERT-XXL with transfer learning from CommonsenseQA achieved the highest accuracy, noting a 4% improvement over simple fine-tuning. Lastly, the effectiveness of fine-tuning extends to commonsense reasoning Del and Fishel (2022) and programming puzzles Schuster et al. (2021), showcasing its broad applicability across puzzle categories.
4 Datasets, Benchmarks and Tasks
Exploring diverse datasets, benchmarks, and tasks is crucial for evaluating LLMs in puzzle-solving. This section examines datasets within our puzzle taxonomy, encompassing formats, evaluation metrics, and methodologies. Figure 2 provides a detailed summary of datasets utilized across the taxonomy’s categories, organized according to puzzle type. The analysis demonstrates LLMs’ versatility and the impact of techniques discussed in §3.
4.1 Rule-based Puzzles
We explore rule-based puzzles to assess LLMs’ understanding within structured, closed-world environments. This includes deterministic puzzles such as Sudoku, Rubik’s Cube, Crosswords, and the 8-puzzle, where solutions follow a set of defined rules. In contrast, stochastic games e.g. Minesweeper, card and social deduction games present variable outcomes from the same actions due to hidden factors. Research predominantly focuses on deterministic puzzles, highlighting a gap in addressing stochastic puzzle uncertainties—a promising direction for future research.
4.1.1 Deterministic Puzzles
Sudoku serves as a prime benchmark for LLMs due to its logical complexity. Noever and Burdick (2021) fine-tune GPT-2 Radford et al. (2019) on 1M Sudoku games, experimenting with compact single-string format, with empty cells represented by "-", and posited that a matrix representation may enhance the model’s learning efficacy. Long (2023) uses nested lists for puzzle representation111e.g. [[3,*,*,2], [1,*,3,*],[*,1,*,3],[4,*,*,1]], finding the Tree-of-Thought (ToT) method most effective, especially for smaller puzzles. Ishay et al. (2023) explore neuro-symbolic approaches across Sudoku, Jobs puzzles and logic puzzles, demonstrating that well-prompted LLMs can accurately generate answer set programming rules.
For Rubik’s Cube and Maze solvers, Noever and Burdick (2021) assess GPT-2’s spatial reasoning using over 2,400 Rubik’s Cube samples and 10K mazes. Despite limited fine-tuning and token constrains, GPT-2 successfully solved the Rubik’s Cube in 1 out of 7 attempts, showing potential despite a high rate of valid though incorrect solutions. Ding et al. (2023) apply multiple methods such as CoT, Self-Consistency, and various Thoughts (ToT, GoT, XoT) on a 2×2×2 Rubik’s Cube using GPT-3.5 and GPT-4. XoT with self-revision emerges as most accurate, significantly outperforming others with a 77.6% success rate.
Exploring LLM versatility, Ding et al. (2023) evaluate the effectiveness of XoT on the spatial 8-Puzzle and numerical Game of 24. The 8-Puzzle’s goal configuration challenges are solved with a remarkable 93.2% accuracy across 419 puzzles using XoT with revision, showcasing superior efficiency over few-shot prompting and CoT. This high accuracy, coupled with a reduced number of LLM invocations, underscores the efficiency and potential of XoT in complex puzzle-solving contexts.
As for Crosswords, Rozner et al. (2021) and Efrat et al. (2021) fine-tune T5 models Raffel et al. (2019) on extensive datasets of individual cryptic clues, revealing T5’s advantage over traditional methods and highlighting areas for improvement, particularly with quick clues and specified answer lengths. Kulshreshtha et al. (2022)’s comparison of BART Lewis et al. (2019) and T5 indicate a sub-30% accuracy for clue-answer tasks, with retrieval-augmented generation transformers surpassing fine-tuned LLMs. Additionally, Yao et al. (2023) apply 5-shot prompting and ToT to GPT-4 on Crossword puzzles significantly improving performance by solving 4 out of 20 puzzles and achieving a 60% word-level success rate.
Feng et al. (2023b) fine-tune two models, "ChessGPT" and "ChessCLIP," using a collection of 3.2M chess puzzles from the Lichess dataset [^2] https://lichess.org/. Each puzzle in the dataset include annotations for its rating, theme, and solution.
At last, Kazemi et al. (2023) unveil BoardgameQA, a dataset featuring multi-choice questions against a backdrop of contradictory facts and rules. Models should navigate through these complexities to provide free-text answers. Their evaluation reveals that fine-tuning BERT-large and T5-XXL with proofs emerges as the most effective method, contrary to few-shot prompting on PaLM with CoT. Moreover, the presence of extra or conflicting information decreases accuracy.
4.1.2 Stochastic Puzzles
The BoardgameQA benchmark Kazemi et al. (2023) also explores scenarios with missing information, which fall under the stochastic puzzle category. It is shown that as missing information increases, the accuracy of fine-tuned models decreases. However, this heightened difficulty does not similarly impact the performance of prompt-tuned and few-shot learning methods, which is likely due to the larger models that were applied.
Minesweeper, known for its hidden information and unpredictability, exemplifies stochastic puzzles, requiring players to deduce mine locations from numerical clues, challenging spatial reasoning. Li et al. (2023) evaluated LLMs on Minesweeper, comparing table and coordinate representations. Even though GPT-3.5 displayed initial understanding, enhancements like few-shot prompting had minimal effects. Conversely, GPT-4 improved mine identification but struggled to complete boards, highlighting Minesweeper’s role in evaluating LLMs’ strategic thinking. Experiments favored the coordinate representation over the table format for aiding LLM comprehension.
Card games, notably Poker, exemplify stochastic puzzles where strategic skill is crucial. Simplified Poker variants require players to infer opponents’ cards and calculate odds amidst hidden intentions. Gupta (2023) found that in Poker’s pre-flop round, ChatGPT and GPT-4 grasp advanced strategies but do not reach Game Theory Optimal (GTO) play. ChatGPT leans towards a conservative approach, while GPT-4 exhibits more aggressive gameplay. Huang et al. (2024) leverage a Reinforcement Learning-trained OPT-1.3B model on all Poker phases revealing superior outcomes in win rates and efficiency, ultimately showcasing LLMs’ adeptness at complex strategies in stochastic settings. An agent that leverages GPT-4 Guo et al. (2023) also achieves significant results in various imperfect information card games.
Social deduction games, including Werewolf and Avalon, blend logical reasoning with complex social dynamics, making them part of the broader stochastic puzzle domain. Such games challenge players to deduce roles involving unpredictable human behavior. Xu et al. (2023b) propose a Werewolf framework using LLMs without tuning, leveraging historical interactions for strategic decisions and showcasing the models’ ability in this context. Similarly, frameworks for Avalon Wang et al. (2023b); Lan et al. (2023) show how LLMs can navigate scenarios demanding social manipulation and deduction, underscoring LLMs’ proficiency in managing the complex interplay of logic and social interaction inherent in such games.
4.2 Rule-less Puzzles
This subsection delves into the diverse datasets related to rule-less puzzles, a category that predominantly encompasses riddles, programming puzzles, and commonsense reasoning challenges. Notably, we specifically focus on puzzles in their traditional sense, thereby excluding code generation datasets, which represent a distinct task type. A majority of rule-less puzzles are structured in a multiple-choice question-answering (QA) format, offering a standardized approach for evaluating LLMs’ inferential reasoning. Benchmarks deviating from this format are specially mentioned, providing a broader perspective on the variety of rule-less puzzle datasets and their implications for LLM performance.
4.2.1 Riddles
RiddleSense Lin et al. (2021) offers a collection of 5.7K vertical thinking riddles, testing pre-trained LMs such as BERT, RoBERTa, ALBERT, and text-to-text QA models including UnifiedQA Khashabi et al. (2020) and T5. Larger LMs generally demonstrate better performance, with UnifiedQA using T5-3B leading, yet struggling with metaphors and counterfactual situations.
Complementing this, BrainTeaser Jiang et al. (2023) introduces 1119 lateral thinking puzzles. It contrasts instruction-based models (ChatGPT, T0, and FlanT5 Chung et al. (2022)) with commonsense ones (including RoBERTa variants and CAR Wang et al. (2023c)). ChatGPT excels in both sentence-based and word-based puzzles, indicating its strength in lateral thinking. However, overall, LLMs still face challenges in exhibiting lateral thinking, with common errors in memorization and commonsense association. This dataset highlights the varied dimensions of reasoning that riddles can test, from vertical logic to lateral inference.
BiRdQA Zhang and Wan (2021) explores the multilingual aspect of riddles, encompassing English and Chinese puzzles, while evaluating monolingual LMs (BERT, RoBERTa), as well as multilingual ones (mBERT, XLM-R Conneau et al. (2019)). The use of brief riddle introductions and hints is also tested. Findings reveal a significant performance gap between LMs and human-level understanding, with monolingual models generally outperforming multilingual ones. Interestingly, additional context such as Wikipedia introductions and hints varied in effectiveness, with such aids benefiting English but not Chinese riddles.
CC-Riddle centers on 27K Chinese character riddles, involving multiple-choice, generative, and retrieval-based formats Xu et al. (2022). Evaluation demonstrates that models encountered difficulties in comprehension and exhibited misunderstandings, revealing the complexities inherent in character-based riddles.
In contrast, PUZZLEQA Zhao and Anderson (2023) offers 558 word puzzles in multiple choice and free-text formats. Larger models, e.g. GPT-3/3.5 show higher accuracy, especially in multiple-choice settings. However, methods such as CoT combined with summarization do not significantly enhance performance, pointing to the ongoing challenges in free-response puzzle solving.
Finally, MARB Tong et al. (2023) encompasses a variety of riddle tasks. Several methodologies including zero-shot, CoT, IEP, and few-shot prompting are tested on models such as GPT-4 and PaLM2-540B Anil et al. (2023). The combination of IEP and CoT emerged as the most effective method, highlighting the value of integrating multiple approaches for diverse riddle types. The dataset also includes commonsense puzzles (§4.2.3), showing similar trends with riddles.
4.2.2 Programming Puzzles
P3 (Python Programming Puzzles) Schuster et al. (2021) offers a range of Python programming challenges, from straightforward string manipulations to complex tasks, such as the Tower of Hanoi and algorithmic puzzles, requiring from the model to find an input that makes the program $f$ return "True". Models applied to these puzzles include enumerative solvers for building Abstract Syntax Trees and autoregressive Language Model Solvers such as GPT-3 and Codex Chen et al. (2021), employing varied prompting techniques. The evaluation metric pass@k, indicates the models’ ability to solve a puzzle within a given number of attempts Chen et al. (2021). Results show a correlation between puzzle difficulty for both models and humans, with descriptive prompts enhancing model performance. Interestingly, models proficient in code completion solved more puzzles with fewer tries, highlighting the importance of specialized capabilities in programming challenges.
Savelka et al. (2023) introduce a dataset comprised of 530 code snippets from programming courses, presenting puzzles in a multiple-choice format. The distinction between questions with and without code snippets offers a unique perspective on LLMs’ problem-solving strategies. The dataset categorizes questions into six types, including true/false and output prediction. GPT models were evaluated, revealing that code inclusion significantly increases puzzle complexity. Accuracy rates vary, with higher performance on completion-oriented questions, suggesting that LLMs’ effectiveness can depend heavily on question format and content.
While both P3 and Programming Snippets Dataset address programming puzzles, they do so in markedly different ways. P3’s focus on finding correct Python program inputs contrasts with the multiple-choice format of the Programming Snippets Dataset. However, both datasets reveal key insights: descriptive prompts aid problem-solving, and question format significantly influences LLM performance.
4.2.3 Commonsense Reasoning Puzzles
True Detective Del and Fishel (2022) presents detective puzzles in long-form stories, challenging LLMs such as GPT-3.5/4 to draw conclusions. Various methods, including CoT and Golden-CoT, are applied, revealing difficulties in making final inferences despite all necessary information being available. Golden-CoT provides the model with the reasoning behind the correct answer, so the model only needs to understand this reasoning and extract the answer. While Vanilla and CoT approaches perform close to random, Golden-CoT demonstrates significantly better accuracy, particularly with GPT-4. However, even with Golden-CoT, GPT-3.5 achieves a solve rate of only 63%, whereas GPT-4 matches human solver results (without access to the reasoning behind the answer).
DetectBench Gu et al. (2023) containing 1200 questions, also evaluates informal reasoning in real-life contexts. It tests methods such as use of hints, various CoT approaches and detective thinking on models including GPT-4, GPT-3.5, GLM-4 and Llama2. Hints emerges as a powerful aid, with larger models generally outperforming smaller ones. The effectiveness of different approaches vary, with detective thinking effectively assisting most of the models.
Both datasets highlight the complexity of real-life reasoning and detective-style puzzles, demonstrating that hints play a crucial role in aiding both human and model performance.
LatEval Huang et al. (2023b) introduces a conversational format with English and Chinese stories, requiring players to ask yes/no questions before providing an answer. GPT-3.5, GPT-4, and various other Chat models are evaluated on their ability to ask relevant questions and maintain consistency with the truth. Larger models do not necessarily show advanced performance in question relevance. However, GPT-4 demonstrates the highest answer consistency, though there is still significant room for improvement. The dataset emphasizes the importance of interactive and conversational reasoning in commonsense understanding.
PuzzTe Szomiu and Groza (2021), with its array of comparison, knights and knaves, and zebra puzzles, represents a potentially rich resource for LLM testing. Despite not yet being applied to LLMs, its generated puzzle answers by Mace4 model finder and Prover9 theorem prover [^3] https://www.cs.unm.edu/ mccune/prover9/ indicate its potential for future LLM evaluations.
The datasets under investigation demonstrate a variety of methods for evaluating commonsense reasoning in LLMs, ranging from detective-style puzzles to interactive story solving. Although larger models generally exhibit better performance, the complexity of these tasks poses significant challenges. Techniques such as sharing additional information through hints show effectiveness in improving outcomes, yet there remains a considerable gap between the performance of models and humans. It is important to note that in this work, we specifically focus on puzzle-oriented benchmarks, excluding general commonsense reasoning datasets e.g. CommonsenseQA, PIQA Bisk et al. (2019) or StrategyQA Geva et al. (2021).
5 Discussion and Future Directions
Applied Methods and Dataset Gaps: Across our puzzle taxonomy, methods such as few-shot prompting, CoT, use of introductions and fine-tuning are commonly employed across most categories. Rule-based deterministic and rule-less commonsense puzzles show the greatest methodological variety, while riddles are also see diverse approaches. In contrast, rule-based stochastic and rule-less programming puzzles exhibit less variety, likely due to fewer studies in these areas. Figure 2 reveals that a substantial number of datasets are available for rule-based deterministic puzzles, such as Sudoku and Rubik’s Cube, as well as a variety of rule-less riddles. This indicates a strong research interest and resource availability in these domains. However, there appears to be a scarcity of datasets for rule-based stochastic puzzles and rule-less programming puzzles. This gap highlights an opportunity for further research and dataset creation to introduce more diverse challenges for advancing the problem-solving capabilities of LLMs. The lack of benchmarks for stochastic puzzles led us to include tasks like card and social deduction games, which share core characteristics with traditional puzzles involving incomplete information. Additionally, neuro-symbolic techniques that translate natural language into code remain notably underutilized in puzzle benchmarks, suggesting a potential area for future exploration.
Comparatively, prompting methods like CoT and ToT enhance complex reasoning abilities without altering the underlying model parameters, yet their effectiveness on smaller LLMs requires further exploration. Moreover, methods such as ToT or GoT necessitate a greater number of model invocations compared to CoT and XoT, which could impact their efficiency and scalability Ding et al. (2023). Fine-tuning, while fundamentally enhances reasoning by altering model parameters, is constrained by its specificity to particular tasks. For instance, models fine-tuned on CSQA have been observed to perform worst on the RiddleSense dataset than models fine-tuned directly on RiddleSense Lin et al. (2021), and a model fine-tuned on Poker games may not perform well on tasks like solving a Rubik’s cube. Finally, puzzle translation methods primarily test the model’s ability to interpret and rephrase problems rather than directly solve them, which tends to have a more limited impact on complex reasoning compared to the other two categories of methods.
Performance Analysis:
Rule-based / Deterministic: Methods such as ToT and XoT (§ 3), typically enhance model reasoning abilities as the complexity of the structure increases Ding et al. (2023). Yet, studies in BoardgameQA and crossword puzzles show generally poor model performance.
Rule-based/Stochastic: Fine-tuning is prevalent here, enabling LLMs to grasp basic rules and simpler scenarios. However, they falter in complex settings that require extensive multi-step reasoning Li et al. (2023).
Rule-less/Riddles & Commonsense: There is a notable performance gap between LLMs and human levels, with methods like CoT improving accuracy but still not matching human evaluation outcomes.
Rule-less/Programming: LLMs find programming puzzles challenging, paralleling human difficulties Schuster et al. (2021). Tasks involving code analysis and reasoning in multiple-choice formats prove particularly tough Savelka et al. (2023).
Furthermore, the format of questions significantly affects puzzle-solving effectiveness. Multiple-choice setups simplify tasks for LLMs by narrowing the solution search space, while free-text formats increase the difficulty level.
Puzzle Generation research is currently limited, likely because the ability to understand and solve puzzles is a prerequisite for generating them. In our survey, we primarily focused on puzzle-solving. The few works we found in puzzle generation reveal mixed results. For instance, GPT-3.5’s attempts to generate puzzles with answers showed poor outcomes Zhao and Anderson (2023). Conversely, the introduction of ACES, an autotelic generation method for diverse programming puzzles, demonstrates how semantic descriptors produced by LLMs can be leveraged for creative puzzle creation Pourcel et al. (2023). Lastly, there are recent works that have studied the generation of crossword puzzles of different languages, utilizing LLMs Zugarini et al. (2024); Zeinalipour et al. (2023b, a).
6 Conclusion
In this survey, we propose a taxonomy of puzzles for evaluating LLMs, categorizing them into rule-based (deterministic and stochastic) and rule-less puzzles (riddles, programming, and commonsense reasoning puzzles). We explore a spectrum of methods for LLM-based puzzle solving, ranging from prompting techniques to neuro-symbolic strategies and fine-tuning. By collating existing datasets in this domain, we provide a comprehensive overview of the resources available for such evaluations. Our analysis identifies current challenges, revealing a difficulty of most methods to successfully solve puzzles, while we outline future directions, emphasizing the need for advanced methodologies and diverse datasets to enhance LLMs’ proficiency in puzzle solving.
Limitations
In this study, we provide a survey of puzzle solving using reasoning of Large Language Models. Despite our best efforts, there may be still some limitations that remain in this paper. Firstly, due to the rapidly evolving nature of this field, we continuously add related approaches and analyses, but it is possible that some recent developments may not be included. Also, due to page constraints, we cannot extensively present all the methods nor provide all the technical details. This might limit the depth of understanding for some readers. Our review only includes methods within 4 years, primarily from sources such as ACL, EMNLP, NAACL, NeurIPS, ICLR, and arXiv. We plan to continue following these sources and adding new methods and datasets. Additionally, all our conclusions §6 are based on empirical analysis. While this provides robust evidence, it may not capture all aspects of the problem. Lastly, as with any survey, our interpretations and conclusions §5 are influenced by our own perspectives and understanding of the field. Other researchers might interpret the same studies differently. Despite these limitations, we believe this study provides a valuable overview of the current state of puzzle-solving using reasoning of Large Language Models.
Appendix A Appendix
A.1 Prompting Topologies
The chain-of-thought (CoT) paradigm involves step-wise explanatory reasoning chains, bolstering capabilities even in zero-shot settings with instructions such as "Let’s think step-by-step" Wei et al. (2022); Kojima et al. (2022). Complementing this, self-consistency generates multiple solution paths, selecting the most coherent one Wang et al. (2022).
Exploring automated feedback, Pan et al. (2023b) examined self-correction within LLMs, noting its varied impact on logical reasoning. While instances of performance enhancement exist Weng et al. (2022); Madaan et al. (2023), broader gains are often elusive, with some strategies even detracting from overall reasoning accuracy Huang et al. (2023a). However, Tyen et al. (2023) highlight the potential of backtracking methods, which, when informed about the specific location of errors, significantly boost the model’s correction abilities.
The Tree-of-Uncertain-Thought (TouT) prompting method structures problem-solving into a tree where each branch explores different uncertain reasoning pathways, allowing for multiple potential solutions Mo and Xin (2023). In contrast, the Tree-of-Thought(s)(ToT) method Yao et al. (2023); Long (2023) focuses on a more linear and deterministic approach, systematically breaking down problems into a single coherent pathway towards a solution. The Graph-of-Thought(s) (GoT) method Besta et al. (2023); Lei et al. (2023) structures problem-solving by mapping out various interconnected reasoning pathways, allowing language models to explore and evaluate multiple solutions simultaneously within a flexible, network-like framework.
A.2 Conventional Methods
AI and Machine Learning methods have long been applied to puzzles and games, with algorithms like Deep Blue Campbell et al. (2002) and AlphaZero Silver et al. (2017) for Chess and Go, renowned for their exceptional results. This section contrasts “traditional” methods used to solve various puzzles with those derived from large language models (LLMs). Note that the aim of this paper isn’t to determine the superior method for each puzzle, but to highlight the distinctive reasoning abilities of LLMs within diverse puzzle contexts. We particularly focus on rule-based puzzles, extensively addressed using conventional methods due to their structured, well-defined environments which require systematic strategies to achieve a solution. Conversely, rule-less puzzles such as riddles primarily test the logical, commonsense reasoning and creativity of models, without a clear path of steps to follow in order to find the solution, so we do not analyze this category.
Chi and Lange (2013) utilized three techniques to solve Sudoku: backtracking, simulated annealing, and alternating projections. The backtracking method, a brute-force depth-first search, consistently resolves puzzles across all difficulty levels, albeit slowly. Constraint programming transforms Sudoku into a constraint satisfaction problem, swiftly enforcing constraints to deduce solutions, often within milliseconds Simonis (2005). These methods always find a solution for Sudoku puzzle, in contrast with LLMs that have not achieved results better than 80% for 5x5 puzzles Long (2023).
In their study on Rubik’s Cube, Chen (2022) employed several traditional methods including Korf’s algorithm Korf (1997), which combines Iterative-Deepening Depth-First Search (IDDFS) with the A* algorithm and a heuristic search database. Both Thistlethwaite’s  [^4] https://www.jaapsch.net/puzzles/thistle.htm and Kociemba’s  [^5] https://kociemba.org/ algorithms utilize group theory and similar search techniques to streamline the solving process, with Kociemba’s version enhancing efficiency by simplifying the group structure. While all these algorithms effectively solve the Rubik’s Cube—a task challenging for LLMs—Korf’s method is particularly noted for its efficiency. Additionally, the study explored a machine learning strategy that integrates Monte-Carlo Tree Search (MCTS) with breadth-first search, yielding more optimized solutions, albeit at a lower efficiency. There have also been various attemts to solve Rubik’s Cube using Reinforcement Learning (RL) like DeepCubeA McAleer et al. (2018); Agostinelli et al. (2019) and others Takano (2023), which although find a solution in relatively few steps are time-consuming, with duration varying from 38.7 to 75.6 seconds Takano (2023).
Mazes are puzzles that can be solved by applying simple algorithms like depth-first search, A* or Trémaux’s algorithm. However these problems are good for testing the spatial reasoning of LLMs. RL has also been utilized to solve mazes with Barj and Sautory (2024) leveraging LLM feedback during training.
In Ding et al. (2023) MCTS has been used to solve Game of 24, 8-Puzzle and Pocket Cube, achieving surpassing many LLM techniques, including CoT, CoT-SC, ToT and GoT. Additionally, Rozner et al. (2021) besides fine-tuning T5 for solving cryptic crosswords, have also used non-neural baselines including a WordNet-based heuristic model, a K-Nearest Neighbours bag of words model and a rule-based model, showing that the fine-tuning of T5 had the best results among them.
Finally, Studholme (2001) proposed a method for solving Minesweeper by considering it as a constraint satisfaction problem (CSP). The core strategy involves transforming the game’s challenges into a set of logical constraints that must be satisfied to avoid mines effectively.
In conclusion, most conventional methods used to solve rule-based puzzles employ deterministic approaches that reliably produce solutions, in stark contrast to the unpredictable nature of LLMs. Another advantage of these traditional methods is their explainability and interpretability, crucial attributes for thoroughly evaluating algorithms and understanding their decision-making processes. However, as demonstrated in the study by Takano (2023), these methods can sometimes exhibit increased time complexity, indicating a potential trade-off between reliability and efficiency.
A.3 Tables
Table 1 delineates the various methods leveraged for puzzle-solving based on the datasets we have collected, illustrating the landscape of current LLM research in this domain. It particularly highlights the extensive methods applied to rule-based deterministic and rule-less commonsense puzzles. The absence of neuro-symbolic techniques and selection inference prompting indicates potential areas for expansion, especially considering their prospective benefits for LLMs grounded in logical reasoning datasets.
Methods Rule-based Puzzles Rule-less Puzzles
Deterministic Stochastic Riddles Programming Commonsense
Prompting - - - - -
Few-shot
Chain-of-Thought
Self-refine
Auto-CoT
Complexity CoT
Plan & Solve
Detective Thinking
Self-Consistency
Tree-of-Thoughts
Tree-of-uncertain-Thoughts
Inferential Exclusion Prompting
Graph-of-Thoughts
Everything-of-thoughts
Hints
Introduction/Summarization
Puzzle Translation - - - - -
Logic
Code
Fine-Tuning
Table 1: Methods used by each category of our taxonomy based on the puzzle benchmarks we collected
The table further reflects the adaptability of certain methods like Chain-of-Thought, few-shot learning and fine-tuning, which are utilized across multiple puzzle types, hinting at their effectiveness. Based on this information, we not only catalogue the current state of method applications in puzzle-solving with LLMs but also highlight opportunities for innovative research in areas yet to be explored.